OpenStudy (anonymous):

using the points (–1, 1) and (0, 6), which of the following is the equation for the line? a. 5y = x – 6 b. y = 5x – 6 c. –5x + y = 6 d. x + 5y = 6

OpenStudy (anonymous):

idk how :/ could you pls explain

OpenStudy (anonymous):

It'll help to first put all the answers in the slope-intercept form: y = mx +b :) Then find the slope using the two points and the point-slope formula

OpenStudy (anonymous):

Well, all you really have to do is make all the answers choices look like B!

OpenStudy (anonymous):

thank you:)

OpenStudy (anonymous):

Then find the slope of the two points: y - y1 = m(x -x1)

OpenStudy (anonymous):

Your welcome, hope this helped :)